OpenStudy (anonymous):

Find the area of a rectangle with length of 6 inches and width of 2 feet???? :/

OpenStudy (anonymous):

1 foot squared

OpenStudy (anonymous):

Ur bomb.

OpenStudy (anonymous):

u gotta both length and width

OpenStudy (anonymous):

2ft=24inches 24*6=96 inches ^2

OpenStudy (anonymous):

u want Area ft or inches?

OpenStudy (anonymous):

if u want area ft than A=96/12ft A=8ft

OpenStudy (anonymous):

i need both ft and in

OpenStudy (anonymous):

A=8ft 24*6=96 inches ^2

OpenStudy (anonymous):

A = 1 sq foot or 144 sq inches